# Dynamic Attention-Guided Diffusion for Image Super-Resolution

### v3

@@ 6 Limitations & Future Work
A notable constraint of this study is its dependence on a good saliency estimation. Even though DINO is trained to be a generic vision encoder, it has known limitations that will reflect on the quality of YODA [11]. Meanwhile, the modularity of YODA allows for the saliency model to be switched once a better one becomes available. Additionally, DINO is explicitly trained on input image resolutions such as $224\times 224$, which may not suffice for image SR applications with much larger spatial sizes of the LR image. An ideal solution would be a scale-invariant extraction of attention maps, e.g., a more extended version of our SIFT-adapted approach. Also, we use a linear correspondence between saliency and the number of steps. We posit that non-linear mappings may affect the quality of results. Lastly, YODA has to be extended for other recent state-of-the-art diffusion-based methods, e.g., latent-based methods like LDM [37] or unsupervised methods based on singular value decomposition like DDRM [24] or DDNM [45], which is orthogonal to our work (see supplementary materials for more details).
@@ Acknowledgment
Complexity of YODA
@@
Identifying Key Regions. To avoid the computational burden of on-the-fly generation, we pre-compute the attention maps prior to training. Table 4 shows the parameter count and throughput of different DINO backbones.
Time-Dependent Masking and Guided Diffusion Process. The integration of attention masks within the diffusion framework introduces minimal computational overhead, thanks to the inherently parallelizable nature of element-wise multiplication and addition, as demonstrated in Equation 9 and Equation 12. Consequently, the predominant factor influencing the overall computational complexity remains the choice of diffusion model, whether it be SR3, SRDiff, or another variant.
@@
This examination of YODA’s complexity highlights its efficiency and the strategic decisions made to balance performance with computational demands. The potential for future improvements underscores YODA’s adaptability and its capacity for integration with emerging sparse diffusion methodologies.
Pipeline of YODA with DINO
@@
More Visual Results on Face SR ($16\times 16\rightarrow 128\times 128$)
@@
Intermediate Results with YODA
Figure 10: Intermediate results of the YODA’s guided diffusion process on CelebA-HQ.
Other State-Of-The-Art Diffusion Models
Method IS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Image Regression [38] 121.1 27.9 0.801
SR3 [38] 180.1 26.4 0.762
LDM-4 (100 steps) [37] 166.3 24.4$\pm$3.8 0.69$\pm$0.14
LDM-4 (big, 100 steps) [37] 174.9 24.7$\pm$4.1 0.71$\pm$0.15
LDM-4 (50 steps, guiding) [37] 153.7 25.8$\pm$3.7 0.74$\pm$0.12
Latent Diffusion Models. Despite the significant advancements brought by Latent Diffusion Models (LDMs) [37], their efficacy in the realm of image SR competes closely with that of SR3 [38], as shown in Table 5. Unfortunately, recent research in this direction focused primarily on text-to-image tasks [17], which makes further comparisons with image SR methods challenging, e.g., SDXL [36], MultiDiffusion [5], or DemoFusion [15]. Nevertheless, their potential for image SR is undeniable. Concerning YODA, we also see great research avenues in combination with LDMs. A critical prerequisite for this synergy is the conversion of attention maps from pixel to latent representations. This aspect has to be investigated in more detail in future work.
[…]
More Details on DINO
DINO [6] is a self-supervised learning approach. involving a teacher and student network. While both networks share the same architecture, their parameters differ. The student network is optimized to match the teacher’s output via cross-entropy loss. During training, both receive two random views of the same input image: the teacher is trained on global views, i.e., $224\times 224$ crops, while the student receives local views, i.e., $96\times 96$ crops. This setup encourages the student to learn “local-to-global” correspondences. In other words, by predicting the teacher’s output, the student learns to infer global information from local views. To prevent mode collapse due to identical architectures, the teacher’s parameters are updated as a moving average of the student’s parameters.
More DINO Attention Maps for YODA
Input
Att.-Head 0
Att.-Head 1
Att.-Head 2
Att.-Head 3
Att.-Head 4
Att.-Head 5
MAX
Figure 11: Comparison of attention maps derived from different DINO layers.

### v4

@@ 6 Limitations & Future Work
A notable constraint of this study is its dependence on a good saliency estimation. Even though DINO is trained to be a generic vision encoder, it has known limitations that will reflect on the quality of YODA [12]. Additionally, DINO is explicitly trained on input image resolutions such as $224\times 224$, which may not suffice for image SR applications with much larger spatial sizes of the LR image. Meanwhile, the modularity of YODA allows for the saliency model to be switched once a better one becomes available. An ideal solution would be a scale-invariant extraction of attention maps, e.g., a more extended version of our SIFT-adapted approach. Lastly, YODA can be extended for other diffusion-based methods, e.g., latent-based methods like LDM [45] or unsupervised methods based on singular value decomposition like DDRM [26] or DDNM [55], which is orthogonal to our work (see appendix for more details).
7 Societal Impact
YODA can significantly benefit fields like medical imaging and remote sensing by improving visual quality. Yet, high-quality SR can also be exploited maliciously by adding realism to deceptive or misleading media content. Using SR methods responsibly and promoting transparency and ethical guidelines in deployment is crucial. Also, the reliance on pre-trained models for attention maps, such as DINO, may introduce biases inherent in the training data. These biases could affect the quality and fairness of the SR results, particularly in diverse or underrepresented populations. Future work should aim to mitigate these biases.
@@ Acknowledgment
In the main text, we presented YODA as an approach to dynamically focus the diffusion process to essential areas in the image. The supplementary material hereby gives further information and visualizations on YODA, such as a discussion on related work, details on DINO, training details, and complexity analysis. It supports understanding the main concepts and ideas examined in the main text.
Related Work
In this section, we discuss other diffusion models that can be applied to YODA and content-aware SR methods that also focus on image content to optimize certain properties in the SR pipeline.
Other State-Of-The-Art Diffusion Models
Method IS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Image Regression [46] 121.1 27.9 0.801
SR3 [46] 180.1 26.4 0.762
LDM-4 (100 steps) [45] 166.3 24.4$\pm$3.8 0.69$\pm$0.14
LDM-4 (big, 100 steps) [45] 174.9 24.7$\pm$4.1 0.71$\pm$0.15
LDM-4 (50 steps, guiding) [45] 153.7 25.8$\pm$3.7 0.74$\pm$0.12
Latent Diffusion Models. Despite the significant advancements brought by Latent Diffusion Models (LDMs) [45], their efficacy in the realm of image SR competes closely with that of SR3 [46], as shown in Table 4. Unfortunately, recent research in this direction focused primarily on text-to-image tasks [19], which makes further comparisons with image SR methods challenging, e.g., SDXL [43], MultiDiffusion [5], or DemoFusion [17]. Nevertheless, their potential for image SR is undeniable. Concerning YODA, we also see great research avenues in combination with LDMs. A critical prerequisite for this synergy is the conversion of attention maps from pixel to latent representations. This aspect has to be investigated in more detail in future work. Our preliminary StableSR (10 epochs in stage 2) results show that our method improves LPIPS from 0.1242 to 0.1239.
Alternative Corruption Spaces. Applying YODA with alternative corruption spaces (not pure Gaussian noise), such as used in InDI [13], I2SB [32], CCDF [11], or ColdDiffusion [4] is also an interesting future research direction. Although our primary focus has been refining and enhancing the diffusion process of specific models through attention-guided masks, we acknowledge the orthogonal potential these approaches represent within the broader context of SR.
Other Content-Aware SR Methods
One category of approaches that draw similarities to YODA is dataset pruning for image SR. Commonly, image SR methods are trained on sub-images cropped from higher-resolution counterparts, such as those found in DIV2K [1]. The central premise behind dataset pruning strategies is the observation that not all sub-images contribute equally to training efficacy. These approaches employ content-aware detectors to prune training data based on metrics like loss values selectively [39] or Sobel norms [14].
A related strategy, ClassSR [51], categorizes different image regions into three levels of reconstruction difficulty - easy, medium, and hard. They propose training specialized models for each category. Related to ClassSR are RAISR [44], SFTGAN [53], RL-Restore [59], and PathRestore [60]. RAISR [44] assigns image patches into clusters and employs a tailored filter to each cluster, utilizing an efficient hashing technique to streamline the clustering process. SFTGAN [53] introduces a spatial feature transform layer that embeds high-level semantic priors, enabling nuanced processing of different image regions with different parameters (i.e., different models). Similarly, RL-Restore [59] and PathRestore [60] divide images into sub-images and employ reinforcement learning to determine the optimal processing pathway for each sub-image. Unlike YODA, which focuses on refining specific image regions with one model, the presented works aim to optimize training datasets by reduction or categorization (thereby employing multiple models tailored to varied reconstruction difficulties).
Similarly, the Multiple-in-One Image Restoration (MiO IR) strategy introduces a novel approach to handling diverse image restoration tasks within a single model [27]. MiO IR employs sequential learning, which allows the model to learn different tasks incrementally and optimize for diverse objectives. Additionally, it utilizes prompt learning - both explicit and adaptive - to guide the model in adapting to various tasks dynamically. While YODA focuses on refining specific image regions, MiO IR’s versatility across tasks offers a potential avenue for expanding YODA’s application to more generalized scenarios.
SkipDiff [37] presents another content-aware SR approach in the context of diffusion models. This method operates through two primary phases: a coarse skip approximation and a fine skip refinement. SkipDiff constructs a preliminary high-resolution image approximation in the first phase in a single step. In the second phase, this image is refined using the classical diffusion pipeline with an adaptive noise schedule. For this, they employ a schedule driven by the characteristics of the input image. Reinforcement learning is integral to this process, as it is trained to find optimal diffusion steps for this phase. This adaptability enables SkipDiff to tailor diffusion to the entire content of an image, contrasting with YODA’s targeted refinement of specific regions. Future research might explore the potential synergies between YODA and SkipDiff, combining their strengths to further enhance content-specific image SR.
Another recent approach in lightweight SR is the Self-Feature Learning (SFL) method proposed by Xiao et al. [57]. This method introduces a locally adaptive involution technique that reduces computational costs by dynamically generating convolutional kernels based on local image content. The SFL model achieves a remarkable trade-off between performance and model complexity by avoiding inter-channel redundancy, making it particularly suitable for resource-constrained devices. Unlike YODA, which focuses on targeted region refinement, SFL employs a dual-path residual module to ensure efficient feature extraction across the entire image, potentially complementing YODA’s targeted strategy.
Details: DINO
DINO [7] is a self-supervised learning approach. involving a teacher and student network. While both networks share the same architecture, their parameters differ. The student network is optimized to match the teacher’s output via cross-entropy loss. During training, both receive two random views of the same input image: the teacher is trained on global views, i.e., $224\times 224$ crops, while the student receives local views, i.e., $96\times 96$ crops. This setup encourages the student to learn “local-to-global” correspondences. In other words, by predicting the teacher’s output, the student learns to infer global information from local views. To prevent mode collapse, the teacher’s parameters are updated as a moving average of the student’s parameters.
Details: Training Parameters
For SR3, we adopted the AdamW [33] optimizer, using a weight decay of 0.0001 and a learning rate of 5e-5. The number of sampling steps is set to $T_{\text{train}}=500$. The number of sampling steps is set to $T_{\text{eval}}=200$. Concerning the denoising architecture, our approach aligns with the SR3 model [46], but we employed residual blocks [22] proposed by Ho et al. instead of those used in BigGAN [23, 6]. Specifically, the configuration includes three ResBlocks, an initial channel size of 64, and a channel multiplier array of [1, 2, 4, 8, 8]. We also employed a NormGroup with a size of 32. For SRDiff, we extracted $40\times 40$ sub-images with a batch size of 16, AdamW [33], a channel size of 64 with channel multipliers [1, 2, 2, 4] and $T=100$.
Details: Complexity of YODA
@@
Identifying Key Regions. To avoid the computational burden of on-the-fly generation, we pre-compute the attention maps prior to training. Table 5 shows the parameter count and throughput of different DINO backbones. While YODA introduces additional complexity for setting up the attention maps, the overhead is minimal. For instance, generating all attention maps for our face SR experiments (i.e., 120,000 images) needed less than two minutes.
Time-Dependent Masking and Guided Diffusion Process. The integration of attention masks within the diffusion framework introduces minimal computational overhead, thanks to the inherently parallelizable nature of element-wise multiplication and addition, as demonstrated in the methodology (see time-dependent masking). Consequently, the predominant factor influencing the overall computational complexity remains the choice of diffusion model, whether it be SR3, SRDiff, or another variant.
@@
This examination of YODA’s complexity highlights its efficiency and the strategic decisions made to balance performance with computational demands.
Details: Analysis Across Attention Regions
This section explains how we created the figure associated with the subsection Analysis across attention regions. The used attention maps are derived from DINO+ResNet. We apply the max aggregation explained in the main paper to create a single attention map per image. The aggregated map is then divided into several attention-value ranges (bins/intervals). Specifically, the attention values range from 0 to 1 (see methodology), and we divide this range into small bins with a step size of 0.01. For each bin, we analyze the regions of the image where the attention values fall within the bin range (e.g., 0.01-0.02, 0.02-0.03, etc.). This allows us to observe how different attention levels correlate with the LPIPS scores and how each SR model performs in these non-overlapping and separated regions.
For each attention value bin, we calculate the LPIPS score between the reference HR image and the output images generated by the different SR techniques (SR3, SR+YODA, and LR/bicubic upsampling). This is done as follows:
  • Mask Creation: For each attention bin, we create a binary mask where pixels in the attention map that fall within the bin are set to 1, and all others are set to 0.
  • Region Extraction: Using this binary mask, we extract the corresponding regions from the HR and SR images.
  • LPIPS Calculation: We compute the LPIPS score between the HR image’s masked regions and the generated images’ masked regions. This process is repeated for each bin across the entire attention range. Masked-out regions do not influence LPIPS because LPIPS is a distance measure of features and masked-out regions in HR, and any SR image leads to an LPIPS of 0.
We calculate the mean LPIPS error within attention bins across the images. A polynomial curve (degree 3) is fitted to the LPIPS scores for each model across the attention value ranges to visualize trends more clearly. This allows us to smooth out potential noise and outliers in the data and observe how each model’s performance changes as we move from high to low-attention regions. Key observations:
  • Higher LPIPS in High Attention Regions: Bicubic upsampling performs poorly in high-attention regions, as indicated by the high LPIPS values. This shows that detail-rich and perceptually important regions are indeed reflected by attention values, as simple upsampling cannot adequately capture the fine details.
  • YODA’s Improvement: YODA shows the most significant improvement in high-attention regions, confirming that YODA refines detail-rich regions more effectively. Moreover, it also shows that YODA is improving every region, thereby truly super-resolving every pixel instead of just replacing the LR and leaving it unchanged.
As DINO+ResNet starts to refine the whole image without dynamic masking at around 0.6, the figure stops there (as it investigated all separated regions by then).
More Visualizations
In the remaining part of the supplementary material, we provide additional visualizations, such as more on attention maps derived from different DINO heads (Figure 7), user study (Figure 10), an overview of the working pipeline of YODA (Figure 9), error maps (Figure 11), comparisons on $16\rightarrow 128$ (Figure 8 and Figure 12), zoomed-in comparisons (Figure 13), intermediate results (Figure 14) as well as intermediate binary masking (Figure 15).
Input
Att.-Head 0
Att.-Head 1
Att.-Head 2
Att.-Head 3
Att.-Head 4
Att.-Head 5
MAX
Figure 7: Comparison of attention maps derived from different DINO heads.
LR
$\uparrow$ PSNR:
$\uparrow$ SSIM:
$\downarrow$ LPIPS:
SR3
23.061
0.6208
0.0504
SR3+YODA
23.289
0.6334
0.0502
HR
Figure 8: SR3 and SR3+YODA reconstructions, $16\rightarrow 128$ (8x). The color shift in SR3 can still be observed (e.g., see background). YODA produces higher-quality images without color shift issues.
@@
Figure 10: Results of our user study. We randomly selected 50 images from the CelebA-HQ dataset and let 45 participants decide which SR predictions are preferred with respect to the given LR image. The scaling of the tested SR prediction task was $16\times 16\rightarrow 128\times 128$ (8x scaling). As a result, YODA+SR3 was preferred in 55.7% of all cases.
Figure 11: Error Maps of SR3 with and without YODA. The brighter, the higher the error. The attention maps generated by DINO are shown above. YODA produces smaller errors, especially in important areas such as face details and hair.
@@
Figure 14: Intermediate results of the YODA’s guided diffusion process on CelebA-HQ (Time steps 189, 168, 147, 126, 105, 84, 63, 42, 21, 0 from left to right).
Figure 15: Intermediate binary masks of YODA’s guided diffusion process on CelebA-HQ.